OpenStudy (anonymous):

Write the equation of the line that passes through (3,5) and (6,6) in slope-intercept form.

OpenStudy (anonymous):

slope is 1/3; y-5=1/3(x-3)

OpenStudy (anonymous):

The Work.?

OpenStudy (anonymous):

then solve for y=

OpenStudy (anonymous):

slope formula: y2-y1/x2-x1 and then point slope formula y-y1=m(x-x1)

OpenStudy (anonymous):

Okay Wait..

OpenStudy (anonymous):

I know the answer is y=1/3x+4 . But how do I do the work.?

OpenStudy (anonymous):

I gave you the formulas. Plug and chug

OpenStudy (anonymous):

Okay This is all I got.

OpenStudy (anonymous):

y2-y1 = 6-5=1 x2-x1 = 6-(3) y=mx+b 6=1/3 (6) + b

OpenStudy (anonymous):

Then, I don't know what to do.

OpenStudy (anonymous):

6=2b so b=3 is what it looks like

OpenStudy (anonymous):

But then the answer won't be 1/3x+4 though. :\

OpenStudy (anonymous):

Yes, I noticed.slope is 1/3; y-5=1/3(x-3) then y-5=1/3x-1 then add 5 to -1 and that's where you get 4

OpenStudy (anonymous):

Can You Show Me The Whole Work.? :\

OpenStudy (anonymous):

I did. substitute the variables in the point slope formula. step 1. then distribute; step 2. Then add 5 to both sides and you are done

OpenStudy (anonymous):

Oh.I think I get it I'll try it out.Thanks.

OpenStudy (anonymous):

But Add 5 To What Both Sides.? IDon't Know What To Do After 6=1/3 (6) +b

OpenStudy (anonymous):

I don't know where you are getting your equation from? y=mx+b??? but something is wrong with it. Look at my work and you add 5 to both sides of the equation.

OpenStudy (anonymous):

First, You're Supposed To Find The Slope. Then You're Supposed To Choose Either Point And Substitute It Along With The Slope Into The Slope Intercept Form. Last, Solve For B. Right.?

OpenStudy (anonymous):

That's The Steps I Followed Like My Teacher Said.

OpenStudy (anonymous):

The Only Thing I'm Confused About Is, Where In The World Does 4 Come From.? >.<